The marginal propensity to consume is 0.7.

How would an initial spending of $1200 affect the GDP?
Answer #1

Answer

The effect will be higher by the multiplier times

Multiplier =1/(1-MPC)

=1/(1-0.7).

=1/0.3

=1/(3/10)

=10/3

change in GDP=change in spending *multiplier

=1200*(10/3)

=$4000

the GDP will increase a maximum of $4000